OpenStudy (anonymous):

Can anyone help me with 8th grade math?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

yes I will see if I can

OpenStudy (anonymous):

4. (07.02 MC) What is the value of x in the equation 4(2x + 14) = 0? (1 point) 9 7 –7 –9 5. (07.02 MC) An equation is shown below: 4(2x – 3) = 1 Which of the following correctly shows the first two steps to solve this equation? (1 point) Step 1: 6x – 7 = 1; Step 2: 6x = 8 Step 1: 6x + 1 = 1; Step 2: 6x = 0 Step 1: 8x – 12 = 1; Step 2: 8x = 13 Step 1: 8x – 3 = 1; Step 2: 8x = 4 6. (07.03 LC) How many solutions does the equation 4y – 4y – 12 = 14 – 2 have? (1 point) One Zero Two Many 7. (07.03 LC) The steps below show the incomplete solution to find the value of x for the equation 8x – 3x – 1 = –2 + 16: Step 1: 8x – 3x – 1 = –2 + 16 Step 2: 8x – 3x – 1 = 14 Step 3: 5x – 1 = 14 Which of these is most likely the next step? (1 point) 5x = 14 5x = 10 5x = 13 5x = 15 8. (07.03 MC) What is the value of y in the equation 2(3y + 7 + 5) = 196 – 16? (1 point) 13 14 26 28 9. (07.03 MC) What is the solution to the following equation? 5(2x – 14) + 23 = 7x – 14 (1 point) 11 14 30 36 10. (07.05 LC) What is the value of a in the equation a = 2 + 3a + 6? (1 point) –8 –4 4 8 11. (07.05 LC) How many solutions does the equation 3x + 6 = –1 – 3 + 4x have? (1 point) Two None One Infinitely many 12. (07.05 MC) Which of these is a simplified form of the equation 9p + 3 = – p + 7 + 2p + 3p? (1 point) 5p = 4 15p = 10 9 = 4 12 = 13 13. (07.05 MC) What is the solution to the equation 4x + 2(x – 3) = 3x + x – 12? (1 point) –3 –1 1 3 14. (07.05 HC) The work of a student to solve the equation 3(3y – 8) = 10 + 3y is shown below: Step 1: 3(3y – 8) = 10 + 3y Step 2: 6y – 5 = 10 + 3y Step 3: 3y – 5 = 10 Step 4: 3y = 15 Step 5: y = 5 In which step did the student first make an error and what is the correct step? (1 point) Step 2: 9y – 24 = 10 + 3y Step 2: 6y + 11 = 10 + 4y Step 3: 6y – 8 = 3(6 + y) Step 3: 3(3y – 8) = 7 + y 15. (07.06 LC) Which statement is true about the equation 4.5z = 4.5z – 4.25 + 6.25? (1 point) It has no solution. It has one solution. It has two solutions. It has infinitely many solutions. 16. (07.06 MC) What is the solution to the equation ? (1 point) n = –2 n = 4 n = n =

OpenStudy (squirrels):

This is not help lol jesus.. ask 1 question at a time.